OpenStudy (anonymous):

What is the solution to the following equation? 5(2x − 12) + 21 = 2x + 41 5 8 10 72

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

@texaschic101

alones (alones):

I came here first lel ;) 5(2x−12)+21=2x+41 simplify both sides of the equation. 5(2x−12)+21=2x+41 (5)(2x)+(5)(−12)+21=2x+41(Distribute) 10x+−60+21=2x+41 (10x)+(−60+21)=2x+41(Combine Like Terms) 10x+−39=2x+41 10x−39=2x+41 Ubtract 2x from both sides. 10x−39−2x=2x+41−2x 8x−39=41 add 39 to both sides. 8x−39+39=41+39 8x=80 At last divide both sides by 8. 8x/8=80/8 x=10 @77777jeannie77777 i fixed it...